OpenStudy (anonymous):

Which expressions would complete this equation so that it has infinitely many solutions 8 + 2(8x – 6) = Choose exactly two answers that are correct. A.2(4x + 7) B.9x – 10 C.16x – 4 D.4(4x – 1)

OpenStudy (anonymous):

@shinalcantara @chosenmatt

OpenStudy (chosenmatt):

A i think

OpenStudy (shinalcantara):

for it have infinite number of solutions, both the right and left expressions must be equal 8 + 2(8x - 6) = 8 + 16x - 12 = 16x - 4 let's check for the choices A.) 2(4x + 7) = 8x + 14 not ok! B.) 9x - 10 not ok! C.) 16x - 4 ok! D.) 4(4x - 1) = 16x - 4 ok!